Suppose that the current 1-year rate (1-year spot rate) and expected 1-year T-bill rates over the following three years (i.e., years 2, 3, and 4, respectively) are as follows:
             1R1 = 4%, E(2r1) = 5%, E(3r1) = 5.5%, E(4r1) = 5.85%
Using the unbiased expectations theory, calculate the current (long-term) rates for 1-, 2-, 3-, and 4-year-maturity Treasury securities. (Round your answers to 2 decimal places.)
Answer #1
1 Year rate 4.00%
2 Year rate = (1.04*1.05)^(1/2)-1 = 4.50%
3 Year rate = (1.04*1.05*1.055)^(1/3)-1 = 4.83%
4 Year rate = (1.04*1.05*1.055*1.0585)^(1/4)-1 = 5.09%